A car travelling at a speed of 29.8 km hr-1 brakes and travels a distance of 0.02km before it comes to a complete stop. Find the time taken for the car to come to rest



1
Expert's answer
2021-04-12T07:00:10-0400

"u=29.8\space km/h\\v=0\space km/h\\s=0.02\space km"

From Newton's laws of motion,

"\Rightarrow v^2-u^2=2as\\\Rightarrow a=\dfrac{-u^2}{2s}\\\Rightarrow a=-22,201\space km/h^2"


"v=u+at"

"\Rightarrow t=\dfrac{-u}{a}=0.00134\space hr=4.83\space s"

t = 4.83 s